OpenStudy (anonymous):

Hi, I'm new on this sit and I was wondering if you'd help me out with this problem: You have 4 identical charges (Q) of 32 microC, each put on the vertex of a rectangle (measuring 2m X 3m). What is the magnitude of the Electrostatic Force on any of the charges? Thanks a lot !!

OpenStudy (anonymous):

The charge on each should be the same. Let's use Coulomb's Law, which is defined as\[E = {1 \over 4 \pi \epsilon_0} { q \over r^2}\]Let's label the vertexes as such, 1) North West 2) North East 3) South East 4) South West Now, the electric field at vertex (1) due to (2), (3), and (4) is as follows:\[E = {1 \over 4 \pi \epsilon_0} \left [{q_2 \over r_{1/2}} + {q_3 \over r_{1/3}} + {q_4 \over r_{1/4}} \right ]\]

OpenStudy (anonymous):

Thanks !! Very helpful....

OpenStudy (anonymous):

I'm sorry. You were looking for electrostatic force. We can use the following expression to find this value. \[F_{1/2} = {1 \over 4 \pi \epsilon_0} {q_1 q_2 \over r^2} {\bf \hat r_{1/2}}\] Keep in mind that force is a vector, so each vertex will have the same magnitude but different direction.

OpenStudy (anonymous):

Good! I was just about to get it wrong... Thanks for the correction